%merlin.mbs apsrev4-1.bst 2010-07-25 4.21a (PWD, AO, DPC) hacked
%Control: key (0)
%Control: author (8) initials jnrlst
%Control: editor formatted (1) identically to author
%Control: production of article title (-1) disabled
%Control: page (0) single
%Control: year (1) truncated
%Control: production of eprint (0) enabled
\begin{thebibliography}{0}%
\makeatletter
\providecommand \@ifxundefined [1]{%
 \@ifx{#1\undefined}
}%
\providecommand \@ifnum [1]{%
 \ifnum #1\expandafter \@firstoftwo
 \else \expandafter \@secondoftwo
 \fi
}%
\providecommand \@ifx [1]{%
 \ifx #1\expandafter \@firstoftwo
 \else \expandafter \@secondoftwo
 \fi
}%
\providecommand \natexlab [1]{#1}%
\providecommand \enquote  [1]{``#1''}%
\providecommand \bibnamefont  [1]{#1}%
\providecommand \bibfnamefont [1]{#1}%
\providecommand \citenamefont [1]{#1}%
\providecommand \href@noop [0]{\@secondoftwo}%
\providecommand \href [0]{\begingroup \@sanitize@url \@href}%
\providecommand \@href[1]{\@@startlink{#1}\@@href}%
\providecommand \@@href[1]{\endgroup#1\@@endlink}%
\providecommand \@sanitize@url [0]{\catcode `\\12\catcode `\$12\catcode
  `\&12\catcode `\#12\catcode `\^12\catcode `\_12\catcode `\%12\relax}%
\providecommand \@@startlink[1]{}%
\providecommand \@@endlink[0]{}%
\providecommand \url  [0]{\begingroup\@sanitize@url \@url }%
\providecommand \@url [1]{\endgroup\@href {#1}{\urlprefix }}%
\providecommand \urlprefix  [0]{URL }%
\providecommand \Eprint [0]{\href }%
\providecommand \doibase [0]{http://dx.doi.org/}%
\providecommand \selectlanguage [0]{\@gobble}%
\providecommand \bibinfo  [0]{\@secondoftwo}%
\providecommand \bibfield  [0]{\@secondoftwo}%
\providecommand \translation [1]{[#1]}%
\providecommand \BibitemOpen [0]{}%
\providecommand \bibitemStop [0]{}%
\providecommand \bibitemNoStop [0]{.\EOS\space}%
\providecommand \EOS [0]{\spacefactor3000\relax}%
\providecommand \BibitemShut  [1]{\csname bibitem#1\endcsname}%
\let\auto@bib@innerbib\@empty
%</preamble>
\end{thebibliography}%


\begin{thebibliography}{99}
\bibitem{Fritch} H. Fritch and P. Minkowski, Phys. Lett. {\bf 62B}, 72 (1976), Eq. 5.
\bibitem{Cabibo} N. Cabibo, Phys. Lett. {\bf 72B}, 333 (1978), Eq. 4.
\bibitem{LSND} C. Athanassopoulos {\it et al.} (LSND Collaboration), Phys. Rev. Lett. {\bf 77}, 3082 (1996).
\bibitem{SuperK} Y. Fukuda, {\it et al.} (Super-Kamiokande Collaboration), Phys. Rev. Lett. {\bf 81}, 1562 (1998). 
\bibitem{KARMEN} B. Armbruster, {\it et al.} (KARMEN Collaboration), Phys. Rev. {\bf D65}, 112011 (2002). 
\bibitem{Kamland} K. Eguchi, {\it et al.} (Kamland Collaboration), Phys. Rev. Lett. {\bf 90}, 021802 (2003). 
\bibitem{Minos} D. G. Michael, {\it et al.} (MINOS Collaboration), Phys. Rev. Lett. {\bf 97}, 191801 (2006).
\bibitem{DayaBay} F.P. An, {\it et al.} (Daya Bay Collaboration), Phys. Rev. Lett. {\bf 108} 171803 (2012). 
\bibitem{Reno} J.K. Ahn, {\it et al.} (Reno Collaboration), Phys. Rev. Lett. {\bf 108}, 191802 (2012). 
\bibitem{DoubleChooz} Y. Abe, {\it et al.} (Double Chooz Collaboration), Phys. Rev. Lett. {\bf 108} 131801 (2012). 
\bibitem{MiniBooNE} A.A. Aguilar-Arevalo,  {\it et al.} (MiniBooNE Collaboration), Phys. Rev. Lett. {\bf 110}, 161801 (2013). 
\bibitem{Pontecorvo} S.M. Bilenky and B. Pontecorvio, Phys. Lett. {\bf 61B}, 248 (1976).
\bibitem{PDG}K.A. Olive {\it et al.} (Particle Data Group), Chin. Phys. C, {\bf 38}, 090001 (2014). 
\end{thebibliography}

\documentclass[preprint,prl,aps]{revtex4-1}
\usepackage{amsmath}
\usepackage{harpoon}
\begin{document}
\date{\today}
\title{Correct $\Delta m^2_{ij}$ Dependence for Neutrino Oscillation Formulae}
\author{Randy A. Johnson}
\affiliation{University of Cincinnati, Cincinnati, OH 45221}
\pacs{14.60.Pq}

\begin{abstract}
The time translation operator for neutrino mass states is often taken to be $e^{-iEt/\hbar}$. This is not relativistically invariant.  In kaon mixing, physicists use $e^{-imc^2\tau/\hbar}$ where $\tau$ is the proper time of the kaon state.  The factor $mc^2\tau$ is the rest frame value of the four vector product $p_\mu x^\mu$ which is an invariant quantity.  If $-i p_\mu x^\mu$ is used in neutrino oscillation formulae instead of $-iEt$, the scale of the $\Delta m^2_{ij}$ is reduced by a factor of two.  

\end{abstract}
\maketitle

In the earliest papers on Dirac neutrino mixing \cite{Fritch, Cabibo}, the operator that has used to translate the neutrino mass states forward in time was taken to be $e^{-iEt/\hbar}$.  This has been used ever since ({\it e.g.}, \cite{LSND, SuperK, KARMEN, Kamland, Minos, DayaBay, Reno, DoubleChooz, MiniBooNE}).  In contrast for kaon mixing, the factor that is used to translate the neutral kaon mass states forward is $e^{-i m c^2 \tau / \hbar}$ where $\tau$ is the proper time of the kaon mass state.  The former phase factor is not relativistically invariant and, therefore, must be incorrect while the latter is proportional to the invariant product $p_\mu x^\mu$ evaluated in the particle's rest frame. It is ironic that Pontecorvio and collaborators, in their initial suggestion of Majorana neutrino oscillations ({\it e.g.}, \cite{Pontecorvo}), used the invariant time translation operator (although they never explicitly stated it).  Subsequent authors did not follow their lead.

Evaluating $p_\mu x^\mu$ in the laboratory frame gives:
\begin{align}
p_\mu x^\mu = \lbrace E,\overrightharp{p}c \rbrace \cdot \lbrace c t,\overrightharp{x} \rbrace & = E_{lab} c t_{lab} - p_{lab} c x_{lab} \nonumber\\
& = \left( \frac{E_{lab}}{\beta} - p_{lab} c\right) x_{lab} \nonumber\\
& = \left( {\frac{E^2_{lab}}{p_{lab} c}} - p_{lab} c\right) x_{lab} \nonumber\\
& = \frac{m^2 c^3}{p_{lab}} x_{lab} \nonumber
\end{align}
This gives a time translation operator of $e^{-i m^2 c^4 L/(E_{lab} \hbar c)}$ if the neutrino has traveled a distance $L$ and if its mass is sufficiently small such that $p_{lab} c \approx E_{lab}$.  The phase of the mass state actually changes twice as fast as with the operator usually used.  The corrected phase factor leads to the two neutrino oscillation formula
\begin{align}
P(\nu_\mu \rightarrow \nu_e ) & = \sin^2 2 \theta \sin^2 \left( \frac{\Delta m^2_{1,2}c^3 L} {2 E_{lab} \hbar} \right) \nonumber \\ 
& = \sin^2 2 \theta \sin^2 \left( 2.534 \frac{\Delta m^2_{1,2} L} { E_{lab}} \right) \nonumber
\end{align}
where, in the last equation, $\Delta m^2_{1,2}$ is in $eV^2$, $L$ is in $km$, and $E$ is in $GeV$. A similar change must be made to the many neutrino formulae.  While this correction does not invalidate any of the measurements of neutrino mass-squared differences, it does change the scale of those differences by a factor of two, $\Delta m^2_{actual}=\Delta m^2_{reported}/2$ for all neutrino oscillation measurements({\it e.g.}, \cite{PDG}).

\end{document}